OpenStudy (anonymous):

MEDAL!!!!!!!!!!!!!!!! ALGEBRA 2!!

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

to find the arc length S=r*theta

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

so its 95.08

OpenStudy (lynfran):

s=8*[95*pi/180] s=8*19/36pi

OpenStudy (lynfran):

we had to convert the degree to radians first

OpenStudy (anonymous):

Ok

OpenStudy (lynfran):

so s=(38/9)pi and since pi=3.14 we get (38/9)*3.14

OpenStudy (anonymous):

so???

OpenStudy (lynfran):

so s=13.257 when rounded =13.26

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

can i get help on one more

OpenStudy (lynfran):

sure

OpenStudy (anonymous):

ok

OpenStudy (lynfran):

 it would be a 315 degree angle so...cos theata=\[(\frac{ \sqrt{2} }{ 2})\], sin theta=\[(\frac{ -\sqrt{2} }{ 2 })\]and tan theta= -1

OpenStudy (lynfran):

ok let me see if i can explain this better

OpenStudy (lynfran):

but those are the answers

OpenStudy (decentnabeel):

good work @LynFran

OpenStudy (lynfran):

dont you just have to state the value of sin,cos and tan...

OpenStudy (anonymous):

yes

OpenStudy (lynfran):

good then the value is about with an image...